OpenStudy (anonymous):

Working Together ,Rosa and Dianne can plant 20 flats of pansies in 3 hours .If Rosa work alone, it takes her 8 hours than it takes Dianne to plant all 20 flats. How long does it take Rosa to plant the pansies by herself.

OpenStudy (anonymous):

it takes her 8 more hours? is it multiple choice honey?

OpenStudy (anonymous):

no one likes these word problems, but we can do it

OpenStudy (anonymous):

yes we can! haha @satellite73

OpenStudy (anonymous):

Not a multiple choice ,I dont like word problems

OpenStudy (anonymous):

@Dumbo do the working and show us how you arrived at your answer

OpenStudy (anonymous):

ok give me a sec here

OpenStudy (anonymous):

I got ...It took Rosa 12 Hours

OpenStudy (anonymous):

to plant 20 flats of pansies.

OpenStudy (anonymous):

yeah, that is what i got too, you did it first

OpenStudy (anonymous):

i solved \[\frac{1}{x}+\frac{1}{x-8}=\frac{1}{3}\] and got \[x^2-14x-24=0\] so \[(x-2)(x-12)=0\] making \(x=12\) or \(x=2\)

OpenStudy (anonymous):

you probably did it an easier way

OpenStudy (anonymous):

So we cant chose x=2

OpenStudy (anonymous):

we both arrived on the same answer but i just like to go easiest way

OpenStudy (anonymous):

Thank you all ...

OpenStudy (anonymous):

ah im too late grrr

OpenStudy (anonymous):

Its Dumbo we are all learning here .I appreciate it so much ...

OpenStudy (anonymous):

no problem thank you for helping me

OpenStudy (anonymous):

You are welcome Dumbo

OpenStudy (anonymous):

haha im glad my real name isnt dumbo

OpenStudy (anonymous):

lol if u want u can call me ads

OpenStudy (anonymous):

Hmm Ads what does that mean